OpenStudy (ms_sp):

Will fan and medal: (Inequalities write each problem and solve): 1. Carson's Parking Garage charges $4.00 for the first 3 hour plus $1.50 for each additional half-hour. Ted has only $1.50 for parking. For how long can Ted park his car in this garage? 2. A taxi charges $2.50 plus $0.20 for each 1/4 of a mile. How many miles can you go if you have $12.50? 3. Melinda's first two test grades in Psych were 87 and 94. What does she need to make on the third test to maintain an average of at least 90?

OpenStudy (ms_sp):

@wildemo @Will.H

OpenStudy (ms_sp):

@brooke..help00

OpenStudy (brooke..help00):

@chpatterson You are very good at things like this. Please explain to him/her what to do.

OpenStudy (ms_sp):

@brooke..help00 thank you for who to suggest!

OpenStudy (ms_sp):

Current they are off @brooke..help00 the person you suggested @chpatterson

OpenStudy (brooke..help00):

@SapphireMoon You may also be able to help:)

OpenStudy (asc.bchs):

I really don't know about the first one, sorry. I'll show the other two though, and hopefully I'm not too late A taxi charges $2.50 plus $0.20 for each 1/4 of a mile. How many miles can you go if you have $12.50? Write an equation 12.50 = 0.20x + 2.50 10 = .2x x = 50 You can go 50 1/4 miles. Let's convert to miles. 50/4 = 12 & 2/4 or 12.5 You can go 12.5 miles Melinda's first two test grades in Psych were 87 and 94. What does she need to make on the third test to maintain an average of at least 90? 90 = (87 + 94 + x) / 3 270 = 87 + 94 + x 92 = x She needs to get a 92%

OpenStudy (ms_sp):

Thank you!